OpenStudy (anonymous):

find the following derivatives: 1. d/dx (squareroot of xe^x) and 2. d/dx ( e^x / cos x)

OpenStudy (reemii):

for the first one, apply the rule for the derivative of a product of functions. for the second one, it's the formula of a quotient. These formulas exist, so just apply.

OpenStudy (anonymous):

not really sure what that means....

OpenStudy (anonymous):

im completely lost on this question

OpenStudy (reemii):

\((f(x)\times g(x))' = f'(x)g(x) + f(x)g'(x)\).

OpenStudy (reemii):

in 1., you have the product of the function \(x\) with the function \(e^x\). apply the formula: \((xe^x)' = x' e^x + x(e^x)'\). Do you see that I just appied the formula?

OpenStudy (reemii):

applied* all you need to do after writing applying the formula is to compute the derivatives that need to be computed (easy).

OpenStudy (reemii):

do you have the answer for question 1. ?

OpenStudy (anonymous):

no

OpenStudy (reemii):

I was almost at the end: \(x'e^x = x(e^x)' = 1e^x+x(e^x) = e^x(1+x)\) .

OpenStudy (anonymous):

i just dont get how you got that its gibberish to me. ugh

OpenStudy (reemii):

There is formula (that you must apply here) that explains how to compute the derivative of a product of functions. It is: \((fg)'=f'g+fg'\). Is it ok until here?

OpenStudy (anonymous):

ok

OpenStudy (reemii):

I didn't see the word "squareroot" .. ouch. do you know how to compute : \((\sqrt{x})'\) ?

OpenStudy (anonymous):

multiply the reciprocal?

OpenStudy (reemii):

?

OpenStudy (anonymous):

no i do not

OpenStudy (reemii):

one (more) formula to remember: \((x^\alpha)' = \alpha x^{\alpha-1}\) so: \((\sqrt x)' = (x^{1/2})' = \frac12x^{-1/2}\).

OpenStudy (reemii):

there are formulas for product, quotient, and composition of functions. - product: i wrote it above - quotient: \((\frac{f}{g})' = \frac{gf'-fg'}{g^2}\) - composition: \((f(g(x)))' = f'g(x))g'(x)\).

OpenStudy (reemii):

1. is \(\sqrt{xe^x}\) if I'm not wrong. Here, it's tricky. there is a product of functions, and there is a composition of functions. You must see put some parts in a box "B".. here let's put \(B=xe^x\), and we see that your function is actually \(\sqrt{B}\).

OpenStudy (reemii):

by the formula for compound functions, \((\sqrt{B})' = \frac{1}{2}B^{-1/2} \times B'\). Then the formula for a product of functions helps yo uto compute \(B'\).

OpenStudy (reemii):

- composition \((f(g(x)))′=f′(g(x))g′(x).\) (i forgot one parenthesis "(" )

OpenStudy (reemii):

is it a bit clearer?

OpenStudy (anonymous):

no im just going to have to go back to book and figure it out but thank you so much!

OpenStudy (reemii):

start with products, then go for compound, then, 1.

OpenStudy (anonymous):

one last question if i can... for this problem i think its undefined but not sure...